# Federico Echenique and Farzad Pourbabaee

### v2

@@ Abstract
We provide a quantitative assessment of welfare in the classical model of risk-sharing and exchange under uncertainty. We prove three kinds of results. First, that in an equilibrium allocation, the scope for improving individual welfare by a given margin (an $\varepsilon$-improvement) vanishes as the number of states increases. Second, that the scope for a change in aggregate resources that may be distributed to enhance individual welfare by a given margin also vanishes. Equivalently: in an inefficient allocation, for a given level of resource sub-optimality (as measured by the coefficient of resource under-utilization), the possibilities for enhancing welfare by perturbing aggregate resources decrease exponentially to zero with the number of states. Finally, we consider efficient risk-sharing in standard models of uncertainty aversion with multiple priors, and show that, in an inefficient allocation, certain sets of priors shrink with the size of the state space.
@@ 1 Introduction
We provide a quantitative assessment of individual and collective welfare in models of risk-sharing, when the state space is large. Our model is completely standard: We consider an exchange economy, with uncertainty, populated by a collection of risk-averse agents who trade to share risk. Uncertainty is captured with a finite, but large, state space. The welfare theorems apply, and Walrasian equilibrium allocations are Pareto optimal; but we are interested in a quantitative assessment of welfare in these economies as the number of states grows large.
Our main results are threefold.
First we consider individual welfare in a Walrasian equilibrium allocation. We quantify welfare by the probability that a random perturbation of an agent’s equilibrium consumption would provide a strict utility improvement: specifically, an improvement of “at least $\varepsilon$” over equilibrium consumption; meaning that the perturbation would remain an improvement if we were to shave off a small fraction $\varepsilon>0$ from the agent’s consumption. (Our first result is actually more general, and quantifies the perturbations that could make at least one agent better off by at least $\varepsilon$, but we focus on a simple special case here in the introduction.) Now, it is obvious that an agent’s consumption can be improved whenever their preferences are monotonic, and the probability of an improvement by perturbation is always strictly positive. Our first result says that the probability that a perturbation is an individual $\varepsilon$-improvement converges (exponentially) to zero as the number of states grows.
To understand our first result, consider the situation of an individual agent $i$ in equilibrium. The usual consumption-maximization diagram is reproduced in Figure 1 on the left. In the figure, there are two states of the world ($d=2$), and an agent’s budget set is depicted in gray. The budget is defined by an equilibrium price $p$, and by the agent’s endowment $\omega_{i}$. The picture shows the agent’s equilibrium consumption $f_{i}$. The agent’s indifference curve, tangent to the budget line at $f_{i}$, is depicted as a dashed curve. The upper contour set is the region to the northeast of the indifference curve that contains all consumptions that are preferred to $f_{i}$.
$f(s_{1})$$f(s_{2})$$f_{i}$$\omega_{i}$$p$$f(s_{1})$$f(s_{2})$$f_{i}$$\omega_{i}$$p$
Figure 1: Individual welfare.
We wish to assess the size of the upper contour set at $f_{i}$: what is the scope for improving $i$’s welfare at the equilibrium allocation $f$? By virtue of the monotonicity of $i$’s preference, the upper contour set is, of course, infinite. It contains all the consumption vectors that are larger, state by state, than $f$, and more. So we consider the size of the upper contour set relative to a ball that is centered at $f_{i}$. Specifically, we measure the size of the dotted region on the right panel of Figure 1. Think of the ball as a set of perturbations of $f_{i}$. [^1] The ball drawn in the figure includes negative consumption, but its radius may be taken to be arbitrarily small so as to avoid negative consumption, if so desired. We may also restrict attention to consumption that is unaffordable at equilibrium prices (the region to the northeast of the budget line) without changing the essence of our results. Indeed, in the latter case, we could draw the figure so that the dotted region has probability close to one; still, with large $d$, its probability vanishes. We wish to quantify how often such perturbations would lead to a utility improvement for $i$. Key to our result is that the improvement should be at least by $\varepsilon$; for simplicity the figure only depicts the strictly better bundles (the case of $\varepsilon=0$).
The consumer would be better off with more consumption in every state, but they would also be willing to accept a tradeoff of more consumption in one state than in another. Of course, the tradeoffs would have to be at terms of trade that are more favorable than equilibrium prices, and guided by the curvature of the agent’s utility function. But the figure shows that there is a significant probability that a perturbation would leave the agent strictly better off, and the upper contour contains much more than the bundles that are larger state-by-state. So does the dotted region. Our first theorem says that, when the number of states is $d$, the probability of the dotted region, calculated for $\varepsilon$ improvements, is bounded above by $\mathrm{e}^{-\varepsilon^{2}d/8}$. In particular, it converges rapidly to zero as $d$ grows large.
To put this bound in perspective, note that the fraction of bundles that are larger, state-by-state, than $f_{i}$ also converges exponentially to zero (it is $(1/2^{d})$). So our result says that, despite the potential for utility improvements by trading off more consumption in some states for less consumption in other, the scope for such potential improvements shrinks when the number of states is large. In a sense, the curvature of utility does not matter in high-dimensional trade. Our bound of $\mathrm{e}^{-\varepsilon^{2}d/8}$ holds for any monotonic utility function.
Our second main result concerns collective welfare in a Pareto optimal allocation. The exercise is similar to our first result, but relies on redistributing aggregate consumption so as to make all agents better off. For our second result, we consider an economy without aggregate uncertainty, and fix an allocation that is Pareto optimal. This time we consider a perturbation of aggregate consumption that can be redistributed to make all agents better off by at least $\varepsilon$ — meaning, again, a utility improvement that remains, even after shaving off a fraction $\varepsilon$ of consumption. We show that a perturbation to aggregate consumption that can be redistributed to improve all agents by at least $\varepsilon$ has vanishingly small probability.
Pareto optimality refers to efficiency with a fixed aggregate endowment, or aggregate consumption. One cannot make all agents better off without changing aggregate consumption. Our result says, when the state space is large, the scope for collective welfare improvements by a given margin $\varepsilon$ is limited, even by changing aggregate consumption. Again it is useful to compare our bound with the fraction of aggregate allocations that are larger than the endowment, state-by-state. The fraction larger allocations shrinks exponentially to zero, as does (according to our results), the allocation that afford a collective improvement.
Our second result can be interpreted in light of Debreu’s coefficient of resource utilization (Debreu, 1951). If we consider an allocation that is not Pareto optimal, then Debreu’s coefficient of resource utilization measures the degree of inefficiency inherent in the allocation. We may again study the probability of an improving perturbation to aggregate consumption that makes agents better off (not by $\varepsilon$ this time, simply strictly better off). Our second main result implies that this probability shrinks to zero exponentially as the number of states grows large. If Debreu’s coefficient of resource utilization is $\mathrm{CRU}$, and Pareto inefficiency implies that $\mathrm{CRU}<1$, then the probability is bounded above by $\mathrm{e}^{-(1-\mathrm{CRU})^{2}d/8}$. [^2] Debreu (1951) proposes $\mathrm{CRU}\in(0,1]$ as a measure of inefficiency of an allocation, and shows that $\mathrm{CRU}=1$ if and only if the allocation is Pareto optimal.
@@
Our third main result is not explicitly about welfare, but instead about the attitudes towards uncertainty by agents who engage in efficient risk sharing. We consider the same setting of economic exchange with no aggregate uncertainty, as in our second result; but we strengthen the assumptions on preferences to focus on utilities with multiple priors. Think, to fix ideas, on the max-min expected utility preferences of Gilboa and Schmeidler (1989) and of a setting with two agents. Now we prove that, if the risk sharing agreement between the two agents can be improved in the Pareto sense, then at least one of the two must have a small set of prior beliefs. The stronger is the level of Pareto improvement, the smaller will be the size of the prior beliefs set. This means that at least one of the two agents must, in some sense, be close to being ambiguity neutral.
@@
Our results rely on the study of concentration of measures, and specifically on the implication of an Isoperimetric inequality to the separation of convex sets. Key is the role of $\varepsilon$, or of $\mathrm{CRU}$, in our discussion above. [^3] Given its role in our results, we discuss the magnitude of $\varepsilon$ in Section 4.2. These magnitudes provide a certain “padding,” or quantitative bound, on the degree of separation of the relevant convex sets. The phenomenon of concentration of measure then has strong consequences for the volume of one of the sets being separated, and in our economic models we can determine which of the sets must have small volume (except for our third application, which is more subtle). Indeed the irrelevance of the curvature in agents’ indifference curves, or upper contour sets, is a familiar phenomenon in the literature on isoperimetric inequalities in high dimension.
@@ 2.1 Notations and Conventions
Let $A$ be a subset of a finite-dimensional normed vector space $(\mathbb{R}^{m},\lVert\cdot\rVert)$. The distance of a point $x\in\mathbb{R}^{m}$ from $A$ is defined by $\mathsf{dist}(x,A)\coloneqq\inf_{a\in A}\lVert x-a\rVert$. By default, the norms used in the paper are the Euclidean $\ell_{2}$ norm. When a particular $p$-norm is used, we refer to the distance function by $\mathsf{dist}_{p}$ and the norm by $\lVert\cdot\rVert_{p}$. For the $\ell_{2}$ norm, we ussualy omit the subscript 2.
We represent the Euclidean open ball centered at $c$ and with radius $r$ by $\mathbb{B}(c,r)\coloneqq\{x\in\mathbb{R}_{+}^{d}:\lVert x-c\rVert<r\}$. When the center is omitted, we take it to be the null vector $\bm{0}$. When the radius is omitted, we assume it is $r=1$. Thus $\mathbb{B}$ denotes the standard (open) unit ball. In the same manner, we denote the uniform probability law on $\mathbb{B}(r)$ by $\mathsf{P}^{r}$.
For two subsets $A,B\subseteq\mathbb{R}^{m}$, we define $\mathsf{dist}(A,B)\coloneqq\inf\left\{\lVert a-b\rVert:a\in A,b\in B\right\}$. We refer to the $\delta$-extension of the subset $A$ by $A^{\delta}=\{x:\mathsf{dist}(x,A)<\delta\}$, that coincides with $A+\delta\,\mathbb{B}$.
Given a measurable subset $A\subseteq\mathbb{R}^{m}$, we denote its Euclidean volume by $\mathsf{Vol}(A)$ that is equal to the Lebesgue integral of the indicator function of $A$ relative to the affine hull of $A$. For example, if $A$ is a $(m-1)$-dimensional surface in $\mathbb{R}^{m}$, then $\mathsf{Vol}(A)$ refers to the surface area of $A$, as opposed to its $m$-dimensional volume (which is zero).
@@ 2.2 Preferences and Uncertainty
We consider a setting with uncertainty. [^4] Using the reinterpretation of commodities in Debreu (1959), our results have implications for many other economic environments. For example, to the textbook version of the model which assumes that consumption is in units of physically distinct goods. Let $S$ be a finite set of possible states of the world, and $d$ be the number of states of the world; $d\coloneqq|S|$.
Consequences are evaluated based on their monetary payoff in $\mathbb{R}_{+}$. The set of acts, mappings from $S\to\mathbb{R}_{+}$, is denoted by $\mathbb{R}_{+}^{d}$, and individual acts are denoted by $f$ and $g$. We endow this set with its natural topology.
Let $\succeq$ be a binary relation on $\mathbb{R}_{+}^{d}$. As usual, we denote the strict part of $\succeq$ by $\succ$, and the associated indifference relation by $\sim$. We say that $\succeq$ is a (weakly monotone) preference relation if it satisfies the following properties:
@@
    That is for every $f\in\mathbb{R}_{+}^{d}$, the sets $\{g:g\succeq f\}$ and $\{g:f\succeq g\}$ are closed subsets of $\mathbb{R}^{d}$.
  • (Monotonicity): For all $f,g\in\mathbb{R}_{+}^{d}$ if $f(s)\geq g(s)$ for all $s\in S$, then $f\succeq g$. Furthermore, if $f(s)>g(s)$ for all $s\in S$, then $f\succ g$.
The space of such preference relations on $\mathbb{R}_{+}^{d}$ is denoted by $\mathcal{P}$.
Given a preference $\succeq$ and an act $f$, the set $\{g:g\succeq f\}$ is called the upper contour set of $\succeq$ at $f$. We say that a preference $\succeq$ is convex if its upper contour sets are convex, for all acts $f$. Convexity jointly with the weak order property and continuity imply that the strict upper contour set, denoted by $\mathcal{U}_{\succeq}^{(0)}\coloneqq\{g:g\succ f\}$, is also convex. We refer to the space of convex preferences by $\mathcal{C}\subset\mathcal{P}$.
@@
An approximate notion of upper contour sets is key to our results.
@@
The approximate upper contour set of preference $\succeq$ at the act $f$ is defined by
$\mathcal{U}_{\succeq}^{(\varepsilon)}(f)=\left\{g\in\mathbb{R}_{+}^{d}:(1-\varepsilon)g\succ f\right\}\,.$
By convexity and continuity of preferences, the approximate upper contour set is a convex and open set. Its convexity is of particular importance to us, as our arguments often rely on the hyperplane separation theorem.
Approximate optimality is often expressed by means of the $\varepsilon$-maximization of some numerical objective function — a utility function $u$ representing a preference $\succeq$. Observe that when $\succeq$ is homothetic (which is the case for many preferences used in applications), then $u$ may be taken to be homogenous of degree one. As a consequence, our notion of approximate optimality translates directly into an approximation of utilities. Indeed, we may then write $\mathcal{U}_{\succeq}^{(\varepsilon)}(f)=\left\{g\in\mathbb{R}_{+}^{d}:(1-\varepsilon)u(g)>u(f)\right\}$.

### v4

@@ Abstract
We study efficient risk sharing among risk-averse agents in an economy with a large, finite number of states. Following a random shock to an initial agreement, agents may renegotiate. If they require a minimal utility improvement to accept a new deal, we show the probability of finding a mutually acceptable allocation vanishes exponentially as the state space grows. This holds regardless of agents’ degree of risk aversion. In a two-agent multiple-priors model, we find that the potential for Pareto-improving trade requires that at least one agent’s set of priors has a vanishingly small measure. Our results hinge on the “shape does not matter” message of high-dimensional isoperimetric inequalities.
@@ 1 Introduction
An economy without aggregate risk is populated by a group of risk-averse agents. The agents reach an efficient risk-sharing agreement that involves, for each of them, a state-contingent consumption plan. [^1] Efficient risk sharing has a long history in economics, see e.g. Borch (1962) and Malinvaud (1973). The agreement provides a certain level of welfare. Before executing the agreement, the agents learn of a random perturbation to the economy that induces aggregate, or systemic, uncertainty. Now the group may wish to renegotiate their initial agreement, but will only do so if they can guarantee themselves a minimal welfare improvement compared to what they had before the perturbation. There are transaction costs to renegotiating that must be covered. We ask: how likely is it that the agents can achieve this improvement? Our main result is that, when the number of states is large, the existence of a welfare-improving renegotiation is extremely unlikely. Its probability decreases exponentially in the number of states.
The diagram on the left in Figure 1 illustrates the situation, albeit in the case with two states ($d=2$) and two agents. Aggregate consumption is fixed at 1 (say 1 million dollars) in every state, represented by the vector $\omega=(1,1)$. Aggregate consumption is state-independent, and thus implies no aggregate risk. The square defined by the origin and $\omega=(1,1)$ is an Edgeworth box that describes all the allocations—all risk-sharing agreements—between the two agents. The Pareto optimal allocations are on the solid curve connecting the origin and $\omega$, while $f=(f_{1},f_{2})$ is a particular Pareto optimal allocation (indicated by the tangency of the agents’ indifference curves at the point $f$, with the usual Edgeworth-box convention that one agent’s coordinate system has been rotated $180^{\circ}$ and its origin coincides with $\omega$). For the allocation $f$, we may consider all the aggregate consumptions that can be decentralized among the two agents so as to obtain higher utility “by $\varepsilon$” than in $f$ [^2] An improvement by $\varepsilon$ means a state-contingent consumption $g$ so that $(1-\varepsilon)g$ is strictly preferred to $f_{i}$. If, for example, preferences are homothetic, this translates into a utility improvement of $\varepsilon/(1-\varepsilon)$%. This set of aggregate bundles consists of all the bundles on the dotted region to the north-east of the dashed curve passing through $\omega$ (a curve that, at $\omega$, has the same slope as the common tangent to agents’ indifference curves at $f$).
[…]
Figure 1: Risk sharing with two states and two agents.
The aggregate consumption $(1,1)$ is subject to a random perturbation, an aggregate shock $z$. Suppose that $z$ is symmetrically distributed around $(0,0)$; in fact, uniformly distributed among all $z$ with $\lVert z\rVert\leq r$—for some $r>0$. Our results say that the probability that $z$ lies in the dotted region, and thus that aggregate risk can be decentralized among the agents in ways that improve individual welfare by $\varepsilon$, is bounded above by $\mathrm{e}^{-\varepsilon^{2}d/8r^{2}}$. For example, when $r=1$, $\varepsilon=0.1$, and $d$ is the number of securities traded on NASDAQ, this bound is about 0.67%. The point is that the probability of an agreeable renegotiation shrinks exponentially to zero as the number of states grows.
The result is surprising because it holds regardless of the agents’ attitudes towards risk. For contrast, imagine instead that agents are extremely risk averse. Any contingent consumption is evaluated according to a worst-case scenario, which is determined by the smallest payoff in any given state. The picture on the right in Figure 1 illustrates the situation. Now an aggregate shock $z$ can only be distributed to make all agents better off if $z(s)>0$ in every state $s$, because if there is a state with less aggregate consumption, then some agent must consume less, and therefore be worse off, under the assumption of extreme risk aversion. The probability of such a perturbation is $1/2^{d}$, which also shrinks exponentially to zero.
Comparing the pictures on the left and right of Figure 1 emphasizes the real message of our results. On the left, agents’ limited risk aversion (the curvature of their indifference curves) leaves significant room for a welfare improvement after the shock $z$. When agents are close to risk-neutral, and $d=2$, the probability of a small $\varepsilon$ improvement after the shock becomes arbitrarily close to $1/2$. As the number of states grows large, however, the probability inexorably decreases to zero. We can even assume that the allocation $f$ is supported by an equilibrium price vector $p$, and restrict attention to shocks $z$ with $p\cdot z>0$, the shocks that have some chance of being welfare-improving. Then the probability in the left Edgeworth box (with $d=2$) is arbitrarily close to $1$; while, again, as $d$ grows, it shrinks exponentially to zero.
Our paper presents three results on economies with many states.
The first result (Theorem 1) concerns a single agent’s welfare. Imagine, again, a group of agents and an efficient allocation $f$. Suppose this time that the allocation is supported by some equilibrium prices; it is part of a Walrasian (or Arrow-Debreu) equilibrium. Given a shock $z$, the group looks for one agent $i$ who is willing to absorb the shock, meaning that $f_{i}+z$ provides them with a welfare improvement of at least $\varepsilon$ over $f_{i}$. The number of agents may be very large, and the group just needs one of them to take on $z$. The probability that there exists such an agent decreases, again, exponentially to zero with the number of states.
The second main result (Theorem 2) concerns an aggregate shock, and the ability to renegotiate an efficient risk-sharing agreement after the shock is realized. This is the result we have described using Figure 1. In Proposition 1, we recast this result for allocations that represent inefficient risk sharing. The degree of inefficiency is measured by Debreu’s coefficient of resource utilization.
@@
Our third main result (Theorem 3) is not explicitly about welfare, but instead about the attitudes towards uncertainty by agents who engage in efficient risk sharing. We consider the same setting of economic exchange with no aggregate uncertainty, as in our second result; but we strengthen the assumptions on preferences to focus on utilities with multiple priors. Think, to fix ideas, of the max-min expected utility preferences of Gilboa and Schmeidler (1989). Now we prove that, if the risk sharing agreement between the two agents can be improved in the Pareto sense, then at least one of the two must have a small set of prior beliefs. The stronger the level of Pareto improvement, the smaller the size of the set of prior beliefs. This means that at least one of the two agents must, in some sense, be close to being ambiguity neutral.
@@
The sources of our, perhaps counterintuitive, results are some equally counterintuitive phenomena in high-dimensional probability. Our results rely on the study of concentration of measures, and specifically on the implication of an isoperimetric inequality for the separation of convex sets. We explore the consequences of the “shapes do not matter” message from the literature in probability for economics. Key is the role of $\varepsilon$, or of the coefficient of resource utilization, in our discussion above. These magnitudes provide a certain “padding,” or bound, on the degree of separation of the relevant convex sets. The phenomenon of concentration of measure then has strong consequences for the volume of one of the sets being separated: one of the separated sets must have small volume. In our model, we can usually determine which of the sets must have small volume, and thus obtain the results we have described. Given its importance, we discuss the role and magnitude of $\varepsilon$ in Section 4.2.
@@ 2.1 Notations and Conventions
Let $A$ be a subset of a finite-dimensional normed vector space $(\mathbb{R}^{m},\lVert\cdot\rVert)$. The distance of a point $x\in\mathbb{R}^{m}$ from $A$ is defined by $\mathsf{dist}(x,A)\coloneq\inf_{a\in A}\lVert x-a\rVert$. By default, the norm used in the paper is the Euclidean $\ell_{2}$ norm. When a particular $p$-norm is used, we refer to the distance function by $\mathsf{dist}_{p}$, and the norm by $\lVert\cdot\rVert_{p}$. For the $\ell_{2}$ norm, we usually omit the subscript 2.
We represent the Euclidean open ball centered at $c$ and with radius $r$ by $\mathbb{B}(c,r)\coloneq\{x\in\mathbb{R}^{d}:\lVert x-c\rVert<r\}$. When the center is omitted, we take it to be the null vector $\bm{0}$. When the radius is omitted, we assume it is $r=1$. Thus $\mathbb{B}$ denotes the standard (open) unit ball.
We denote the uniform probability law on $\mathbb{B}(r)$ by $\mathsf{P}^{r}$. We further define the class of (Borel) probability measures supported on $\mathbb{B}(r)$ that are absolutely continuous with respect to $\mathsf{P}^{r}$, with their Radon-Nikodym derivative almost everywhere bounded above by a constant $\kappa$, as follows:
$\mathcal{M}^{r}_{\kappa}\coloneq\left\{\mathsf{P}^{r}_{\kappa}\in\Delta\left(\mathbb{B}(r)\right):\frac{\mathrm{d}\mathsf{P}^{r}_{\kappa}}{\mathrm{d}\mathsf{P}^{r}}(z)\leq\kappa,\,\,\text{a.e.}\,\,z\in\mathbb{B}(r)\right\}\,.$
For example, the $d$-dimensional standard Gaussian measure restricted to the $r$-ball has a bounded Radon-Nikodym derivative with respect to $\mathsf{P}^{r}$, with the constant $\kappa\leq\mathrm{e}^{r^{2}/2}$ (see Section 6.1).
For two subsets $A,B\subseteq\mathbb{R}^{m}$, we define $\mathsf{dist}(A,B)\coloneq\inf\left\{\lVert a-b\rVert:a\in A,b\in B\right\}$. We refer to the $\delta$-extension of the subset $A$ by $A^{\delta}=\{x:\mathsf{dist}(x,A)<\delta\}$, which coincides with $A+\delta\,\mathbb{B}$.
Given a measurable subset $A\subseteq\mathbb{R}^{m}$, we denote its Euclidean volume by $\mathsf{Vol}(A)$, which is equal to the Lebesgue integral of the indicator function of $A$ relative to the affine hull of $A$. For example, if $A$ is a $(m-1)$-dimensional surface in $\mathbb{R}^{m}$, then $\mathsf{Vol}(A)$ refers to the surface area of $A$, as opposed to its $m$-dimensional volume (which is zero).
@@ 2.2 Preferences and Uncertainty
We consider a setting with uncertainty. [^3] Using the reinterpretation of commodities in Debreu (1959), our results have implications for many other economic environments. For example, for the textbook version of the model which assumes that consumption is in units of physically distinct goods. Let $S$ be a finite set of possible states of the world, and $d$ be the number of states of the world; $d\coloneq|S|$.
Consequences are evaluated based on their monetary payoff in $\mathbb{R}$. The set of acts, mappings from $S\to\mathbb{R}$, is denoted by $\mathbb{R}^{d}$, and individual acts are denoted by $f$ and $g$. We endow this set with its natural topology.
Let $\succeq$ be a binary relation on $\mathbb{R}^{d}$. As usual, we denote the strict part of $\succeq$ by $\succ$, and the associated indifference relation by $\sim$. We say that $\succeq$ is a (weakly monotone) preference relation if it satisfies the following properties:
@@
    That is, for every $f\in\mathbb{R}^{d}$, the sets $\{g:g\succeq f\}$ and $\{g:f\succeq g\}$ are closed subsets of $\mathbb{R}^{d}$.
  • (Monotonicity): For all $f,g\in\mathbb{R}^{d}$ if $f(s)\geq g(s)$ for all $s\in S$, then $f\succeq g$. Furthermore, if $f(s)>g(s)$ for all $s\in S$, then $f\succ g$.
The space of such preference relations on $\mathbb{R}^{d}$ is denoted by $\mathcal{P}$. We say that $\succeq$ is a strongly monotonic preference relation if, in addition to the weak order and continuity properties, it satisfies strong monotonicity: Namely, for all $f,g\in\mathbb{R}^{d}$ if $f(s)\geq g(s)$ for all $s\in S$ and $f\neq g$, then $f\succ g$. We denote the space of strongly monotone preferences by $\mathcal{P}^{\text{sm}}\subset\mathcal{P}$.
Given a preference $\succeq$ and an act $f$, the set $\{g:g\succeq f\}$ is called the upper contour set of $\succeq$ at $f$. We say that a preference $\succeq$ is convex if its upper contour sets are convex, for all acts $f$. Convexity jointly with the weak order property and continuity imply that the strict upper contour set, denoted by $\mathcal{U}_{\succeq}^{(0)}\coloneq\{g:g\succ f\}$, is also convex. We refer to the space of convex preferences by $\mathcal{C}\subset\mathcal{P}$. We interpret convexity in our model as a basic property of risk, or uncertainty aversion (Yaari, 1969).
@@
An approximate, or robust, notion of upper contour sets is key to our results.
@@
The $\varepsilon$-approximate upper contour set of the preference $\succeq$ at the act $f$ is defined by
$\mathcal{U}_{\succeq}^{(\varepsilon)}(f)=\left\{g\in\mathbb{R}^{d}:(1-\varepsilon)g\succ f\right\}\,.$
When $\succeq$ is continuous and convex, the approximate upper contour set $\mathcal{U}_{\succeq}^{(\varepsilon)}(f)$ is convex and open. Convexity is often of particular importance to us, as some of our arguments rely on a hyperplane separation theorem.
Approximate optimality is often expressed by means of the $\varepsilon$-maximization of some numerical objective function—a utility function $u$ representing a preference $\succeq$. Observe that when $\succeq$ is homothetic (which is the case for many preferences used in applications), then $u$ may be taken to be homogeneous of degree one. As a consequence, our notion of approximate optimality translates directly into an approximation of utilities. Indeed, we may then write $\mathcal{U}_{\succeq}^{(\varepsilon)}(f)=\left\{g\in\mathbb{R}^{d}:(1-\varepsilon)u(g)>u(f)\right\}$. Then an act $g$ is in $\mathcal{U}_{\succeq}^{(\varepsilon)}(f)$ if and only if it represents an $\varepsilon/(1-\varepsilon)$utility improvement over $f$.